OpenStudy (anonymous):

how many integer values of (a) can x^2+ax+6 be factored? what are they?

OpenStudy (anonymous):

to find possible values of a just find the sum of integers that multiply to six. integers that multiply to 6 3x2 1x6 -1x-6 -2x-3 -1+(-6)=-7 -2+(-3)=-5 3+2=5 1+6=7 four different values of a can be factored

OpenStudy (anonymous):

which list is it?

OpenStudy (anonymous):

is a=2

OpenStudy (anonymous):

what are the 4 values????